OpenStudy (anonymous):

Find the fifth roots of 243(cos 240° + i sin 240°).

OpenStudy (anonymous):

I solved it. Here's some steps for future references. First, write the number in trig form: r^5 = 243 and theta = 240 degrees. Now, I must take the 5th root of 243 which is 3. So, one root is "3". Apply deMoivre's Theorem and divide 240 degrees by 5 and add (k+360/3) with k going from 0 to 2: (240/5)+(0+360/3) = 168 (240/5)+(1+360/3) = 169 (240/5)+(2+360/3) = 170 Your roots are: 3, 168, 169, 170